# Constrained Stochastic Recursive Momentum Successive Convex Approximation

### v3

@@ Appendix E Proof of Corollary 2 and Corollary 3
Under non-adaptive step sizes $\eta_{t}=\eta=\tfrac{{\bar{k}}}{T^{1/3}}$ and $\beta_{t}=\beta=c\eta^{2}$, the modified results in Lemmas 1 and 2 stated as follows,
$\frac{1}{16L^{2}\eta}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t+1}\right\|^{2}\right]\leq\frac{(1-\beta)}{16L^{2}\eta}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]+\frac{\eta}{8}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]+\frac{\beta^{2}}{8L^{2}\eta}{\mathbb{E}}\left[(G_{t+1})^{2}\right],$ (103)
and
@@
${\mathbb{E}}\left[F({\mathbf{x}}_{t+1})-F({\mathbf{x}}_{t})\right]\leq\frac{\eta}{2}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]-\frac{\eta}{4}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right].$ (104)
@@
By introducing a Lyapunov function $\Psi_{t}={\mathbb{E}}\left[F({\mathbf{x}}_{t})-F^{\star}\right]+\tfrac{1}{16L^{2}\eta}{\mathbb{E}}[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}]$ and adding (103) with (104), we obtain
@@
$\Psi_{t+1}-\Psi_{t}$ $\leq\frac{\alpha}{16L^{2}}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]-\frac{\eta}{8}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]+\frac{c^{2}\eta^{3}}{8L^{2}}{\mathbb{E}}\left[G_{t+1}^{2}\right],$ (105)
with $\alpha=8L^{2}\eta+\frac{(1-\beta)}{\eta}-\frac{1}{\eta}$. Taking telescopic sum from $t=1\ldots T$,
$\Psi_{T+1}-\Psi_{1}$ $\leq\frac{\alpha}{16L^{2}}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]-\frac{\eta}{8}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]+\frac{c^{2}\eta^{3}}{8L^{2}}\sum_{t=1}^{T}G^{2}$ (106)
Also, we have that
$\alpha$ $=8L^{2}\eta-\frac{\beta}{\eta}=-2L^{2}\eta\left(\frac{c}{2L^{2}}-4\right)=-2L^{2}d^{\prime}\eta$ (107)
where $d^{\prime}:=\frac{c}{2L^{2}}-4>0$. Here, since ${\mathbb{E}}[\left\|{{\boldsymbol{\varepsilon}}}_{1}\right\|^{2}]\leq\sigma^{2}$ from Assumption A5, we can bound
$\Psi_{1}\leq F({\mathbf{x}}_{1})-F^{\star}+\frac{\sigma^{2}}{16L^{2}\eta_{0}}=B_{1}+\frac{\sigma^{2}}{16L^{2}{\bar{k}}}.$ (108)
Rearranging (105) we can write,
@@
$\tfrac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]\leq\frac{M^{\prime}_{T}}{T^{2/3}}={\mathcal{O}}\left(\frac{1}{T^{2/3}}\right)$ (109)
where $M^{\prime}_{T}:=8\frac{B_{1}}{{\bar{k}}}+\frac{\sigma^{2}}{2L^{2}\bar{k}^{2}}+\frac{c^{2}\bar{k}^{2}}{L^{2}}G^{2}$. Similarly,
@@
$\tfrac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]\leq\frac{\bar{M^{\prime}}_{T}}{d^{\prime}T^{2/3}}={\mathcal{O}}\left(\frac{1}{T^{2/3}}\right)$ (110)
where $d^{\prime}=\tfrac{1}{2L^{2}}(c-8L^{2})$. Finally, using similar inequalities as (41) and (42) we obtain the results of Corollary 2.
@@
Also, under non-adaptive step size, in light of Corollary 2, (60) can be bounded as,
@@
$\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\pi}}_{t}\right\|\right]^{2}$ $\leq\tfrac{2}{T}\sum_{t=1}^{T}\left(2L+\frac{4B_{U}L^{2}}{\rho^{2}}\right)^{2}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|\right]^{2}+\tfrac{2}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\varepsilon}}_{t}\right\|\right]^{2}$
$\leq{\mathcal{O}}(T^{-2/3})$ (111a)
@@
$\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}[{\hat{{\boldsymbol{\lambda}}}}_{t}^{\mathsf{T}}{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})]$ $\geq-\frac{2B_{U}L^{2}}{\rho^{2}}{\mathbb{E}}[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}]={\mathcal{O}}(T^{-2/3}).$ (111b)
@@
$\min_{1\leq t\leq T}({\mathbb{E}}\left[\left\|{\boldsymbol{\pi}}_{t}\right\|\right]^{2}-{\mathbb{E}}[{\hat{{\boldsymbol{\lambda}}}}_{t}^{\mathsf{T}}{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})])$ $\leq\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\pi}}_{t}\right\|\right]^{2}-\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}[{\hat{{\boldsymbol{\lambda}}}}_{t}^{\mathsf{T}}{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})]$
$\leq{\mathcal{O}}(T^{-2/3})$ (111c)
Hence, and we obtain the desired result in Corollary 3

### v4

@@ Appendix E Proof of Corollary 2 and Corollary 3
Under non-adaptive step sizes, specifically $\eta_{t}=\tfrac{{\bar{k}}}{(w+t)^{1/3}}$ and $\beta_{t+1}=c\eta_{t}^{2}$. The results of Lemma 1 and 2 the step sizes $\eta_{t}$ can be taken outside the expectation. Consequently, the modified result in Lemma 1 becomes,
$\frac{1}{16L^{2}\eta_{t}}$ ${\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t+1}\right\|^{2}\right]\leq\frac{1-\beta_{t+1}}{16L^{2}\eta_{t}}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]$
$+\frac{\eta_{t}}{8}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]+\frac{\beta_{t}^{2}}{8L^{2}\eta_{t}}{\mathbb{E}}\left[(G_{t+1})^{2}\right].$ (103)
Lemma 2 is also modified as follows.
@@
${\mathbb{E}}\left[F({\mathbf{x}}_{t+1})-F({\mathbf{x}}_{t})\right]\leq\frac{\eta_{t}}{2}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]-\frac{\eta_{t}}{4}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right].$ (104)
@@
Note that here only Assumption A6 is required instead of Assumption A3 (as $\eta_{t}$ can taken outside the expectation in (88)). Following the proof structure of Theorem 1, and introducing the Lyapunov function $\Psi_{t}={\mathbb{E}}\left[F({\mathbf{x}}_{t})-F^{\star}\right]+\tfrac{1}{16L^{2}\eta_{t-1}}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]$, we obtain,
@@
$\Psi_{t+1}-\Psi_{t}$ $\leq\frac{\alpha_{t}}{16L^{2}}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}\right]-\frac{\eta_{t}}{8}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]+\frac{c^{2}\eta_{t}^{3}}{8L^{2}}{\mathbb{E}}\left[G_{t+1}^{2}\right],$ (105)
where $\alpha_{t}:=8L^{2}\eta_{t}+\frac{1-\beta_{t+1}}{\eta_{t}}-\frac{1}{\eta_{t-1}}$. Just like (35), we can again write
$\frac{{\bar{k}}}{\eta_{t}}-\frac{{\bar{k}}}{\eta_{t-1}}$ $=(w+t)^{1/3}-(w+t-1)^{1/3}$
$\leq\frac{1}{3(w+t-1)^{2/3}}=\frac{\eta_{t-1}^{2}}{3{\bar{k}}^{2}}.$ (106)
Now, similar to (36)-(40), we have
$\alpha_{t}$ $\coloneq 8L^{2}\eta_{t}+\frac{(1-\beta_{t+1})}{\eta_{t}}-\frac{1}{\eta_{t-1}}$
$=8L^{2}\eta_{t}+\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}-\frac{\beta_{t+1}}{\eta_{t}}$ (107)
$\leq 8L^{2}\eta_{t-1}+\frac{\eta_{t-1}^{2}}{3\bar{k}^{3}}-c\eta_{t-1}\leq-2L^{2}d^{\prime}\eta_{t-1}$ (108)
where $d^{\prime}:=\tfrac{1}{2L^{2}}(c-8L^{2}-\tfrac{1}{3\bar{k}^{3}})>0$. Therefore the first term on the right of (105) can be dropped. Taking telescopic sum of (105) from $t=1\ldots T$ and rearranging,
$\frac{1}{8}\sum_{t=1}^{T}\eta_{t}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]$ $\leq\Psi_{1}-\Psi_{T+1}+\frac{c^{2}G^{2}}{8L^{2}}\sum_{t=1}^{T}\eta_{t}^{3}$ (109)
Similar to (43) we can bound,
$\sum\nolimits_{t=1}^{T}\eta_{t}^{3}={\bar{k}}^{3}\sum\nolimits_{t=1}^{T}\frac{1}{w+t}\leq{\bar{k}}^{3}\ln(1+\frac{T}{w})$
$\qquad=\mathcal{O}(\ln(1+\frac{T}{w}))$ (110)
Like (45) we can bound,
$\Psi_{1}\leq F({\mathbf{x}}_{1})-F^{\star}+\frac{\sigma^{2}}{16L^{2}\eta_{0}}=B_{1}+\frac{w^{1/3}\sigma^{2}}{16L^{2}{\bar{k}}}.$ (111)
Hence, finally, using similar inequalities as (42) we lower bound (109),
$\Delta_{T}^{2}\leq M^{\prime}_{T}\frac{(w+T)^{1/3}}{T}$ (112a)
where $M^{\prime}_{T}:=8\frac{B_{1}}{{\bar{k}}}+\frac{w^{1/3}\sigma^{2}}{2L^{2}\bar{k}^{2}}+\frac{c^{2}{\bar{k}}^{2}G^{2}}{L^{2}}\ln(1+\frac{T}{w})$. Similarly, we get
@@
$\tfrac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]\leq\frac{M^{\prime}_{T}(w+T)^{1/3}}{T}={\tilde{\mathcal{O}}}\left(\frac{1}{T^{2/3}}\right)$ (112b)
@@
$\tfrac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|\right]\leq\sqrt{\frac{M^{\prime}_{T}(w+T)^{1/3}}{d^{\prime}T}}={\tilde{\mathcal{O}}}\left(\frac{1}{T^{1/3}}\right)$ (112c)
where $d^{\prime}:=\tfrac{1}{2L^{2}}(c-8L^{2}-\tfrac{1}{3\bar{k}^{3}})>0$. Hence, we obtain the desired result in Corollary 2.
@@
Also, under non-adaptive diminishing step size, in light of Corollary 2, (59) can be bounded as,
$\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\pi}}_{t}\right\|\right]^{2}$
@@
$~~~~\leq\tfrac{2}{T}\sum_{t=1}^{T}\left(2L+\frac{4B_{U}L^{2}}{\rho^{2}}\right)^{2}{\mathbb{E}}\left[\left\|{\boldsymbol{\delta}}_{t}\right\|\right]^{2}+\tfrac{2}{T}\sum_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\varepsilon}}_{t}\right\|\right]^{2}$
$~~~~\leq{\tilde{\mathcal{O}}}(T^{-2/3})$ (113a)
@@
$\tfrac{1}{T}\sum_{t=1}^{T}{\mathbb{E}}[{\hat{{\boldsymbol{\lambda}}}}_{t}^{\mathsf{T}}{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})]\geq-\frac{2B_{U}L^{2}}{\rho^{2}T}\sum_{t=1}^{T}{\mathbb{E}}[\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}]={\tilde{\mathcal{O}}}(T^{-2/3}).$ (113b)
@@
$\min_{1\leq t\leq T}({\mathbb{E}}\left[\left\|{\boldsymbol{\pi}}_{t}\right\|\right]^{2}-{\mathbb{E}}[{\hat{{\boldsymbol{\lambda}}}}_{t}^{\mathsf{T}}{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})])$
$~~~~~~~~~\leq\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}\left[\left\|{\boldsymbol{\pi}}_{t}\right\|\right]^{2}-\tfrac{1}{T}\sum\nolimits_{t=1}^{T}{\mathbb{E}}[{\hat{{\boldsymbol{\lambda}}}}_{t}^{\mathsf{T}}{\mathbf{g}}({{\hat{{\mathbf{x}}}}_{t}})]$
$~~~~~~~~~\leq{\tilde{\mathcal{O}}}(T^{-2/3})$ (113c)
Hence, we obtain the desired result in Corollary 3. ∎